# Centaur: a foundation model of human cognition

### v2

@@ Abstract
Establishing a unified theory of cognition has been a major goal of psychology [1, 2]. While there have been previous attempts to instantiate such theories by building computational models [1, 2], we currently do not have one model that captures the human mind in its entirety. Here we introduce Centaur, a computational model that can predict and simulate human behavior in any experiment expressible in natural language. We derived Centaur by finetuning a state-of-the-art language model on a novel, large-scale data set called Psych-101. Psych-101 reaches an unprecedented scale, covering trial-by-trial data from over 60,000 participants performing over 10,000,000 choices in 160 experiments. Centaur not only captures the behavior of held-out participants better than existing cognitive models, but also generalizes to new cover stories, structural task modifications, and entirely new domains. Furthermore, we find that the model’s internal representations become more aligned with human neural activity after finetuning. Taken together, Centaur is the first real candidate for a unified model of human cognition. We anticipate that it will have a disruptive impact on the cognitive sciences, challenging the existing paradigm for developing computational models.
@@ Introduction
In contrast to this, most contemporary computational models – whether in machine learning or the cognitive sciences – are domain-specific. They are designed to excel at one particular problem and that problem alone. Take, for instance, AlphaGo – a computer system created by Google DeepMind to master the game of Go [9]. Even though the system can play this particular game at an impressive level, it can do not much beyond that. A similar pattern emerges in the cognitive sciences. Prospect theory, one of the most influential accounts of human cognition, for instance, offers valuable insights into how people make choices [10], but it tells us nothing about how we learn, plan, or explore.
Psych-101: 160 psychological experiments, 60,092 individual participants, 10,681,650 human choices, 253,597,411 text tokensabCentaur: a foundation model of human cognition Multi-armed banditsy Decision-making Memory Supervised learning Markov decision processes Miscellaneousy In this task, you have to repeatedly choose between two slot machines labeled B and C. When you select one of the machines, you will win or lose points. Your goal is to choose the slot machines that will give you the most points.
@@
Taken together, these results demonstrate that it is possible to discover domain-general models of human cognition in a data-driven manner. We believe that Centaur is the first real candidate for a unified model of human cognition and that it offers many opportunities to obtain a better understanding of the human mind.
@@ Results
We built Centaur on top of the open-source language model Llama 3.1 70B – a state-of-the-art model pre-trained by Meta AI [13] (hereafter, we refer to this model simply as Llama). Having a large language model as the backbone allowed us to rely on the vast amounts of knowledge that is present in these models [16]. The training process involved finetuning on Psych-101 using a parameter-efficient finetuning technique known as quantized low-rank adaptation (QLoRA) [17]. QLoRA leaves the parameters of the base model intact while adding so-called low-rank adapters, which contain only a few additional, trainable parameters. In our case, we added low-rank adapters of rank eight to all non-embedding layers as illustrated in Figure 1b. With these settings, the newly added parameters amount to 0.15% of the base model’s parameters. We then trained the model for one epoch on the entire data set using a standard cross-entropy loss. We masked out the loss for all tokens that do not correspond to human responses, thereby ensuring that the model focuses on capturing human behavior and not on completing experimental instructions. The entire training process took approximately five days on an A100 80GB GPU. Further details on the finetuning procedure are provided in the Methods section.
@@
Figure 2: Performance on Psych-101. a, Pseudo-R2 values for different models across experiments. A value of zero corresponds to prediction at chance level while a value of one corresponds to perfect predictability of human responses. Missing bars indicate performance below chance level. Centaur outperforms both Llama and a collection of domain-specific cognitive models in almost every experiment. Note that we only included experiments for which we have implemented a domain-specific cognitive model in this graphic and merged different studies using the same paradigm. A full table for all experiments can be found in the Supplementary Information. b, Model simulations on the two-step task. The plot visualizes probability densities over reward and a parameter indicating how model-based learning was for people and simulated runs of Centaur. c, Model simulations on the horizon task. The plot visualizes probability densities over reward and an information bonus parameter for both people and simulated runs of Centaur. d, Model simulations on a grammar judgement task. The plot visualizes probability densities over true and estimated scores (i.e., number of correct responses out of twenty) for both people and simulated runs of Centaur.
@@
We evaluated Centaur on different types of held-out data to demonstrate that it robustly captures human behavior. In our first analysis, we tested whether it can predict behavior of participants that were not part of the training data. For this, we split each transcribed experiment into two parts, and used 90% of participants for training and retained 10% for testing. We measured goodness-of-fit to human choices using a pseudo-R2 measure, which normalizes the log-likelihood of a model by that of a randomly guessing model [18]. In this measure, a value of zero corresponds to prediction at chance level while a value of one indicates perfect predictability. [^1] Note that the (unknown) noise ceiling is in general lower than one, which can only be attained when predicting deterministic behavior. Figure 2a presents the result of this analysis, comparing Centaur against the base model without finetuning and collection of domain-specific models that represent the state-of-the-art in the cognitive science literature. We observed that all models predict human behavior above chance level in the majority of experiments. While there was substantial variance in predictability between the experiments, finetuning always improved goodness-of-fit. The average improvement across experiments after finetuning was $0.14$ (Centaur pseudo-R2 = $0.50$; Llama pseudo-R2 = $0.36$).
@@
Furthermore, we compared Centaur against the previously mentioned collection of domain-specific cognitive models. These models include, amongst others, the generalized context model [19], a prospect theory model [20], and various reinforcement learning models [21, 22]. Further technical details about the modeling can be found in the Supplementary Information. We observed that Centaur outperforms domain-specific cognitive models in all but one experiment. The average improvement in predicting human behavior over the domain-specific cognitive models was $0.18$ (Centaur pseudo-R2 = $0.50$; cognitive models pseudo-R2 = $0.32$).
[…]
@@
The previous analyses have focused on predicting human responses conditioned on previously executed behavior. We may ask whether Centaur can also generate human-like behavior when simulated in an open-loop fashion (i.e., when feeding its own responses back into the model). This setting arguably provides a much stronger test for the model’s capabilities [23]. To check whether Centaur survives this test, we ran open-loop simulations in three different experimental paradigms and inspected the distributions of statistics that resulted from these simulations. The corresponding results can be found in Figure 2b-d. We found that Centaur performs at human-level in all of these simulations, confirming that it can generate meaningful open-loop behavior. Furthermore, Centaur’s distributions are well-aligned with the human population, demonstrating that Centaur does not merely model the behavior of the average participant but rather the distribution over trajectories produced by the entire population. For example, in the two-step task – a well-known paradigm to tease apart model-free and model-based reinforcement learning [21] – Centaur produced trajectories in which learning is purely model-free, purely model-based, and mixtures thereof (see Figure 2b).
@@
Figure 3: Evaluation in different held-out settings. a, Pseudo-R2 values for the two-step task with a modified cover story [24]. b, Pseudo-R2 values for a three-armed bandit experiment [25]. c, Pseudo-R2 values for an experiment probing logical reasoning [26]. Centaur outperforms both Llama and domain-specific cognitive models when faced with modified cover stories, problem structures, and entirely novel domains.
@@
First, we investigated whether Centaur is robust in the face of changes to the cover story. For this analysis, we relied on data collected by Feher da Silva and Hare [24], who conducted a study using the aforementioned two-step task. In addition to the canonical cover story (spaceships traveling to foreign planets in search of treasures), their study introduced a novel cover story involving magical carpets. Importantly, Psych-101 includes experiments using the canonical spaceship cover story [27, 28] but no experiments with the magical carpet cover story. Yet, we still found that Centaur captures human behavior in the magical carpet experiment of Feher da Silva and Hare (see Figure 3a). Like in our previous analysis, we observed an improvement after finetuning, as well as a favorable goodness-of-fit when compared to a domain-specific cognitive model (Centaur pseudo-R2: $0.26$; Llama pseudo-R2: $0.10$; cognitive model pseudo-R2: $0.12$).
@@
In a second out-of-distribution evaluation, we probed whether Centaur is robust to modifications in task structure. To test this, we exposed it to a paradigm known as Maggie’s farm [25]. Maggie’s farm extends the horizon task paradigm – a two-armed bandit task used to detect different types of exploration strategies [22] – by adding a third choice option. Psych-101 encompasses several two-armed bandit experiments (including the horizon task) but not Maggie’s farm or any other three-armed bandit experiments. [^2] It does, however, contain multi-armed bandit experiments with more than three choice options. Thus, this analysis provides a test of Centaur’s robustness to structural task modifications. We found that Centaur captures human behavior on Maggie’s farm as shown in Figure 3b. We again observed a benefit of finetuning, as well as a favorable goodness-of-fit compared to a domain-specific cognitive model, which did not generalize well to this setting (Centaur pseudo-R2: $0.62$; Llama pseudo-R2: $0.43$; cognitive model pseudo-R2: $0.11$).
@@
Finally, we investigated whether Centaur can capture human behavior even in entirely novel domains. In this context, we considered a study investigating logical reasoning [26]. While Psych-101 includes probabilistic and causal reasoning problems, we purposefully excluded any studies involving logical reasoning. Like in the previous analyses, there was again a positive effect of finetuning (Centaur pseudo-R2: $0.18$; Llama pseudo-R2: $0.05$; see Figure 3c). Note that we did not compare to any domain-specific cognitive model in this setting, as it is unclear how to construct a model that would make any meaningful transfer from training data that does not include any related problems.
@@
Taken together, these analyses cement Centaur’s ability to generalize to experiments outside of its training data. We additionally verified the model on a collection of benchmarks from the machine learning literature [29, 30]. We found that Centaur remains stable in performance-based benchmarks, even improving over the base model in some of them [30]. Furthermore, in benchmarks that measure human alignment, we observed a shift towards human-like characteristics. Figure 4a depicts this improved alignment on a low-dimensional embedding derived from ten behavioral metrics in CogBench, a benchmark to test the cognitive abilities of large language models [29]. A more detailed description of these additional benchmarking results can be found in the Supplementary Information.
@@
Despite only being trained to match human behavior, we also wondered whether Centaur’s internal representations become more aligned with human neural activity. To check whether this is the case, we conducted two analyses in which we predicted human neural activity using the model’s internal representations [31, 32]. We first conducted a whole-brain analysis, in which we predicted fMRI measurements of people performing the two-step task [33]. We extracted recordings before each choice and after feedback. We then aggregated human neural activity in each region and regressed the aggregated activity on Centaur’s internal representations. This procedure was then repeated separately for each participant and region. Further details can be found in the Methods section. Figure 4b shows the resulting R2 values across layers for both Centaur and Llama. We found that Centaur’s representations consistently outperform Llama’s representations in predicting human neural activity, suggesting that finetuning a model on large-scale behavioral data aligned its internal representations to human neural activity. It is worth noting that this type of analysis was only possible due to the expressivity of Centaur’s representations, and that using representations of a traditional cognitive model barely predicted human neural activity better than chance (dashed grey line in Figure 4b).
@@
Figure 4: Alignment between humans and Centaur. a, Multidimensional scaling embedding of the ten behavioral metrics in CogBench [29] for different models. b, R2 values indicating how well human neural activity in the two-step task [33] can be decoded using Centaur’s internal representations extracted from different layers. c, R2 values indicating how well human neural activity in a sentence-reading task [34] can be decoded using Centaur’s internal representations extracted from different layers.
We verified these results in a second analysis, for which we relied on a previously collected data set involving fMRI measurements of people reading simple, six-word sentences [34]. We closely followed the protocol of the original study and predicted aggregated neural activity across participants in the language network. We repeated this procedure for representations extracted from different layers in both Centaur and Llama. Predictability peaked at around layer 20 as shown in Figure 4c. This peak is consistent with the hypothesis that the intermediate layers of such models contain the most amount of information [35]. More importantly, however, this analysis confirmed that finetuning consistently improves decodability across all layers. Even though the improvement is somewhat small, this setting provides a particularly strong test for the generalization abilities of Centaur, given that the stimuli used in this study are far away from its training data and that it was never trained to capture neural activity.

### v3

@@ Abstract
Establishing a unified theory of cognition has been a major goal of psychology [1, 2]. While there have been previous attempts to instantiate such theories by building computational models [1, 2], we currently do not have one model that captures the human mind in its entirety. A first step in this direction is to create a model that can predict human behavior in a wide range of settings. Here we introduce Centaur, a computational model that can predict and simulate human behavior in any experiment expressible in natural language. We derived Centaur by finetuning a state-of-the-art language model on a novel, large-scale data set called Psych-101. Psych-101 reaches an unprecedented scale, covering trial-by-trial data from over 60,000 participants performing over 10,000,000 choices in 160 experiments. Centaur not only captures the behavior of held-out participants better than existing cognitive models, but also generalizes to new cover stories, structural task modifications, and entirely new domains. Furthermore, we find that the model’s internal representations become more aligned with human neural activity after finetuning. Taken together, our results demonstrate that it is possible to discover computational models that capture human behavior across a wide range of domains. We believe that such models provide tremendous potential for guiding the development of cognitive theories and present a case study to demonstrate this.
@@ Introduction
In contrast to this, most contemporary computational models – whether in machine learning or the cognitive sciences – are domain-specific. They are designed to excel at one particular problem and that problem alone. Take, for instance, AlphaGo – a computer system created by Google DeepMind to master the game of Go [9]. Even though the system can play this particular game at an impressive level, it can do not much beyond that. A similar pattern can be observed in the cognitive sciences. Prospect theory, one of the most influential accounts of human cognition, for instance, offers valuable insights into how people make choices [10], but it tells us nothing about how we learn, plan, or explore.
Refer to caption160 psychological experiments, 60,092 individual participants, 10,681,650 human choices, 253,597,411 text tokensabCentaur: a foundation model of human cognition Multi-armed banditsy Decision-making Memory Supervised learning Markov decision processes Miscellaneousy In this task, you have to repeatedly choose between two slot machines labeled B and C. When you select one of the machines, you will win or lose points. Your goal is to choose the slot machines that will give you the most points.
@@
Taken together, our results demonstrate that it is possible to discover computational models that capture human behavior across a wide range of domains. We believe that such a predictive model offers many direct opportunities to obtain a better understanding of the human mind [16, 17] and present a case study that demonstrates this potential.
@@ Results
We built Centaur on top of the open-source language model Llama 3.1 70B – a state-of-the-art model pre-trained by Meta AI [13] (hereafter, we refer to this model simply as Llama). Having a large language model as the backbone allowed us to rely on the vast amounts of knowledge that is present in these models [18]. The training process involved finetuning on Psych-101 using a parameter-efficient finetuning technique known as quantized low-rank adaptation (QLoRA) [19]. QLoRA relies on a frozen four-bit quantized language model as a base model. While the parameters of the base model are left unchanged, it adds so-called low-rank adapters, which contain only a few additional, trainable parameters (typically represented in a half-precision floating-point format). In our case, we added low-rank adapters of rank $r=8$ to all non-embedding layers (i.e., all linear layers of the self-attention mechanisms and the feedforward networks) as illustrated in Figure 1b. With these settings, the newly added parameters amount to 0.15% of the base model’s parameters. We then trained the model for one epoch on the entire data set using a standard cross-entropy loss. We masked out the loss for all tokens that do not correspond to human responses, thereby ensuring that the model focuses on capturing human behavior and not on completing experimental instructions. The entire training process took approximately five days on an A100 80GB GPU. Further details on the finetuning procedure are provided in the Methods section.
@@
Figure 2: Goodness-of-fit on Psych-101. a, Difference in log-likelihood of Centaur and Llama relative to a domain-specific cognitive model for each experiment. A value of zero corresponds to the goodness-of-fit of the domain-specific cognitive model while a value above zero indicates improved goodness-of-fit to human responses. Error bars correspond to the standard error of the mean, taken over responses. Centaur outperforms both Llama and a collection of domain-specific cognitive models in almost every experiment. Note that we only included experiments for which we have implemented a domain-specific cognitive model in this graphic and merged different studies using the same paradigm. A full table for all experiments can be found in the Supplementary Information. b, Model simulations on the horizon task. The plot visualizes probability densities over reward and an information bonus parameter for both people and simulated runs of Centaur. c, Model simulations on the two-step task. The plot visualizes probability densities over reward and a parameter indicating how model-based learning was for both people and simulated runs of Centaur. d, Model simulations on a social prediction game. The plot visualizes probability densities over accuracies of predicting human strategies and strategies of an artificial agent with matched statistics for both people and simulated runs of Centaur.
@@
We evaluated Centaur on different types of held-out data to demonstrate that it robustly captures human behavior. In our first analysis, we tested whether it can predict behavior of participants that were not part of the training data. For this, we split each transcribed experiment into two parts, and used 90% of participants for training and retained 10% for testing. We measured goodness-of-fit to human choices using negative log-likelihoods averaged across responses. Figure 2a presents the result of this analysis, comparing Centaur against the base model without finetuning and collection of domain-specific models that represent the state-of-the-art in the cognitive science literature. While there was substantial variance in predictability between the experiments, finetuning always improved goodness-of-fit. The average difference in log-likelihoods across experiments after finetuning was $0.14$ (Centaur negative log-likelihood: $0.44$; Llama negative log-likelihood: $0.58$; $t=-144.22,p\leq 0.0001$).
@@
Furthermore, we compared Centaur against the previously mentioned collection of domain-specific cognitive models. These models include, amongst others, the generalized context model [20], a prospect theory model [21], and various reinforcement learning models [22, 23]. Further technical details can be found in the Supplementary Information. We observed that Centaur outperforms domain-specific cognitive models in all but one experiment. The average difference in predicting human behavior to the domain-specific cognitive models was $0.13$ (cognitive models negative log-likelihood: $0.56$; $t=-178.7,p\leq 0.0001$).
@@
The previous analyses have focused on predicting human responses conditioned on previously executed behavior. We may ask whether Centaur can also generate human-like behavior when simulated in an open-loop fashion (i.e., when feeding its own responses back into the model). This setting arguably provides a much stronger test for the model’s capabilities and is sometimes also referred to as model falsification [24]. To check whether Centaur survives this test, we ran open-loop simulations in three different experimental paradigms and inspected the distributions of statistics that resulted from these simulations. First, we simulated Centaur on the horizon task paradigm, a two-armed bandit task used to detect different types of exploration strategies [23]. We found that Centaur achieved a performance comparable to human participants, as well that it engaged in a similar level of uncertainty-guided, directed exploration (see Figure 2b), a pattern that is absent in many contemporary language models [14]. We also observed that Centaur’s behavior is well-aligned with the human population, demonstrating that it does not merely model the behavior of the average participant but rather the distribution over trajectories produced by the entire population. For example, in the two-step task – a well-known paradigm to tease apart model-free and model-based reinforcement learning [22] – Centaur, like human subjects, produced trajectories in which learning is purely model-free, purely model-based, and mixtures thereof (see Figure 2c). Lastly, we verified that Centaur fails at predicting non-human behavior. For this, we considered a study that required participants to predict either human responses or responses of an artificial agent with matched statistics in four canonical economic games [25]. Mirroring the results of the original human study, Centaur accurately predicted human responses ($64\%$ accuracy) but struggled to predict artificial responses ($35\%$ accuracy; see Figure 2d). Taken together, these results demonstrate that Centaur exhibits human-like characteristics across various settings, confirming that it can generate meaningful open-loop behavior.
@@
Figure 3: Evaluation in different held-out settings. a, Negative log-likelihoods for the two-step task with a modified cover story [26]. b, Negative log-likelihoods for a three-armed bandit experiment [27]. c, Negative log-likelihoods for an experiment probing logical reasoning [28] with items based on the Law School Admission Test (LSAT). Centaur outperforms both Llama and domain-specific cognitive models when faced with modified cover stories, problem structures, and entirely novel domains. Error bars correspond to the standard error of the mean, taken over responses.
@@
First, we investigated whether Centaur is robust in the face of changes to the cover story. For this analysis, we relied on data collected by Feher da Silva and Hare [26], who conducted a study using the aforementioned two-step task. In addition to the canonical cover story (spaceships traveling to foreign planets in search of treasures), their study introduced a novel cover story involving magical carpets. Importantly, Psych-101 includes experiments using the canonical spaceship cover story [29, 30] but no experiments with the magical carpet cover story. Yet, we still found that Centaur captures human behavior in the magical carpet experiment of Feher da Silva and Hare (see Figure 3a). Like in our previous analysis, we observed an improvement after finetuning, as well as a favorable goodness-of-fit when compared to a domain-specific cognitive model (Centaur negative log-likelihood: $0.51$; Llama negative log-likelihood: $0.63$; cognitive model negative log-likelihood: $0.61$; all pairwise comparisons significant with $p\leq 0.0001$).
@@
In a second out-of-distribution evaluation, we probed whether Centaur is robust to modifications in task structure. To test this, we exposed it to a paradigm known as Maggie’s farm [27]. Maggie’s farm extends the horizon task paradigm by adding a third choice option. Psych-101 encompasses several two-armed bandit experiments (including the horizon task) but not Maggie’s farm or any other three-armed bandit experiments. [^1] It does, however, contain multi-armed bandit experiments with more than three choice options. Thus, this analysis provides a test of Centaur’s robustness to structural task modifications. We found that Centaur captures human behavior on Maggie’s farm as shown in Figure 3b. We again observed a benefit of finetuning, as well as a favorable goodness-of-fit compared to a domain-specific cognitive model, which did not generalize well to this setting (Centaur negative log-likelihood: $0.42$; Llama negative log-likelihood: $0.62$; cognitive model negative log-likelihood: $0.98$; all pairwise comparisons significant with $p\leq 0.0001$).
@@
Finally, we investigated whether Centaur can capture human behavior even in entirely novel domains. In this context, we considered a study investigating logical reasoning [28]. While Psych-101 includes probabilistic and causal reasoning problems, we purposefully excluded any studies involving logical reasoning. Like in the previous analyses, there was again a positive effect of finetuning (Centaur negative log-likelihood: $1.65$; Llama negative log-likelihood: $1.92$; $t=-50.39,p\leq 0.0001$; see Figure 3c). Note that we did not compare to any domain-specific cognitive model in this setting, as it is unclear how to construct a model that would make any meaningful transfer from training data that does not include any related problems.
We consolidated these results by analyzing Centaur on six additional out-of-distribution experimental paradigms that were not part of the training data in any shape or form (including moral decision-making [31], economic games [32], naturalistic category and reward learning [33], behavioral propensities [34], and a deep sequential decision task [35]). Centaur robustly captured human behavior in all of these settings, while smaller and non-finetuned models did not do so consistently. We present the corresponding results in the Supplementary Information.
In addition to analyzing human choice data, we also examined whether Centaur can predict human response times. Hick’s law [36] suggests that individual response times are a linear function of the response entropy. Therefore, we extracted nearly 4,000,000 response times for a subset of experiments in Psych-101 and fitted three linear mixed effects models each predicting response time based on the response entropy derived from a different model. We found that the response entropy derived from Centaur captured a larger proportion of the variance in response times (conditional R2: 0.58) than those derived from Llama (conditional R2: 0.4) and the cognitive models (conditional R2: 0.38), thereby highlighting Centaur’s ability to predict measures beyond pure choice data.
@@
To demonstrate that the model does not degrade on problems it was pretrained for, we furthermore verified it on a collection of benchmarks from the machine learning literature [37, 38]. We found that Centaur remains stable in performance-based benchmarks, even improving over the base model in some of them [38]. Finally, in benchmarks that measure human alignment, we observed a shift towards human-like characteristics. Figure 4a depicts this improved alignment on a low-dimensional embedding derived from ten behavioral metrics in CogBench, a benchmark to test the cognitive abilities of large language models [37]. A more detailed description of these additional benchmarking results can be found in the Supplementary Information.
@@
Despite only being trained to match human behavior, we also wondered whether Centaur’s internal representations become more aligned with human neural activity. To check whether this is the case, we conducted two analyses in which we predicted human neural activity using the model’s internal representations [39, 40]. We first conducted a whole-brain analysis, in which we predicted fMRI measurements of people performing the two-step task [41]. For this, we relied on data collected in a previous study [41], involving 94 participants each making 300 choices. Participants were tested on either the magical carpet cover story (that we already used in one of our earlier generalization analyses) or an abstract cover story. Neither of these two cover stories were part of Centaur’s training data. We extracted recordings from models’ residual stream before each choice and after feedback. We then aggregated human neural activity in each region and regressed the aggregated activity on Centaur’s internal representations. This procedure was then repeated separately for each participant and region. Further details can be found in the Methods section. Figure 4b shows the resulting Pearson correlation coefficients across layers for both Centaur and Llama. We found that Centaur’s representations consistently outperform Llama’s representations in predicting human neural activity (all $p\leq 0.001$), suggesting that finetuning a model on large-scale behavioral data aligned its internal representations to human neural activity. It is worth noting that this type of analysis was only possible due to the expressivity of Centaur’s representations, and that using representations of a traditional cognitive model led to a substantially drop in performance (dashed line in Figure 4b).
@@
Figure 4: Human alignment. a, Multidimensional scaling embedding of the ten behavioral metrics in CogBench [37] for different models. b, Pearson correlation coefficients indicating how well human neural activity in the two-step task [41] can be decoded using Centaur’s internal representations extracted from different layers. c, Pearson correlation coefficients indicating how well human neural activity in a sentence-reading task [42] can be decoded using Centaur’s internal representations extracted from different layers. Control refers to a model that uses representations extracted from a randomly-initialized transformer model with matched architecture.
[…]
Model-guided scientific discovery using Psych-101 and Centaur
Psych-101 and Centaur both constitute valuable tools for scientific discovery. In the following section, we present an example for how each of them can be used to improve our understanding of human decision-making. The individual steps of this process are illustrated in Figure 5.
Refer to caption Refer to caption
Expert 1 (90%) 1 0
Expert 2 (80%) 0 1
Expert 3 (70%) 0 1
Expert 4 (60%) 1 1
Which product is superior in quality?Original studyDeepSeek-R1 explanation The participant employed a two-step decision-making strategy. First, they determined which product had the majority of positive ratings (1s) across all experts. If the products were tied in the number of positive ratings, the participant then considered the rating from the highest validity expert to break the tie. Scientific regret minimization Product A Product B Response 1 0 0 1 0 1 1 1 A 0 1 1 0 1 0 0 0 B 0 1 1 0 1 0 0 0 B 1 0 0 0 0 1 1 0 A 0 1 1 1 1 0 0 1 B 1 0 0 1 0 1 1 1 A 0 1 1 0 1 0 0 0 B Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Model-guided scientific discovery. We used Psych-101 and Centaur to guide the development of a cognitive model for a multi-attribute decision-making study. Each panel shows the Akaike information criterion (AIC) for the set of models considered at the given stage, starting with the models considered in the original study. We first asked DeepSeek-R1 to generate an explanation for human responses and formalized the resulting verbal strategy into a formal computational model. We then further refined this model through scientific regret minimization using Centaur as a reference model. Eight data points are visualized for which Centaur makes accurate predictions but the DeepSeek-R1-discovered model does not. We then used this information to design a domain-specific cognitive model that is as predictive as Centaur yet still interpretable.
[…]